OpenStudy (tomektomeknyc):

2 cos ^2 6 is identical to : a) 1- cos 12 b) 1 + cos 12 c) 1+ cos 3 d) 1/2 sin ^2 6

OpenStudy (raden):

cos^2 (x) = (1 + cos(2x))/2 or 2cos^2 (x) = (1 + cos(2x)) now, what will happend if you replace x = 6 to both sides?